# DIAS: A Dataset and Benchmark for Intracranial Artery Segmentation in DSA sequences

### v3

@@ Abstract
Digital subtraction angiography (DSA) is universally acknowledged as the gold standard for examining lesion angioarchitecture, elucidating arterial blood supply dynamics, and guiding endovascular interventions. The automatic segmentation of intracranial arteries (IA) in DSA, which is pivotal for quantifying vascular morphology, plays an essential role in computer-assisted stroke research and clinical practices. Nevertheless, research in this specific domain remains constrained, primarily owing to the unavailability of publicly datasets for IA segmentation within the research community. Currently, the predominant focus of methodologies lies in the segmentation of single-frame DSA using in-house datasets. These methods, limited by the partial inclusion of contrast in single-frame DSA, encounters challenges in rendering a precise representation of vascular structures. In this paper, we introduces DIAS, a dataset specifically developed for IA segmentation in DSA sequences. A comprehensive benchmark has been established for evaluating DIAS, covering fully, weakly, and semi-supervised segmentation methods. Specifically, we propose a vessel sequence segmentation network that captures the spatiotemporal representation of intravascular contrast for segmenting vessels in DSA sequences. For weakly-supervised learning, we propose a novel scribble learning-based image segmentation framework, incorporating both scribble supervision and consistency regularization. Furthermore, we introduce a random patch-based self-training framework that harnesses unlabeled DSA sequences to improve segmentation performance. Our extensive experiments on the DIAS dataset demonstrate the effectiveness of these methods as potential baselines for future research and clinical applications. The dataset and code will be publicly available at https://github.com/lseventeen/DIAS.
@@ I Introduction
Cerebrovascular diseases substantially contribute to global mortality and long-term disability [1]. Recent decades have witnessed a decline in mortality from intracranial artery (IA) diseases like intracranial arterial stenosis (ICAS), middle cerebral artery occlusion (MCAO), and intracranial aneurysm, thanks to advancements in medical and imaging technology [2]. Among these, Digital Subtraction Angiography (DSA), an imaging modality, visualizes the flow of a contrast agent through vessels by capturing a sequence of images over a duration typically spanning 3 to 15 seconds, with a sampling rate between 3 to 5 frames per second [3]. It effectively removes bony structures by subtracting pre-contrast images, thereby enhancing the visibility of vessels filled with the contrast agent. Owing to its inherently superior spatial and temporal resolution, DSA can accurately reveal lesion details in instances where computed tomography angiography (CTA) and magnetic resonance angiography (MRA) may not offer a definitive diagnosis [2]. Thus, DSA is widely recognized as the gold standard for investigating lesion angioarchitecture, deciphering the dynamics of arterial blood supply, and guiding endovascular intervention [4].
@@
Conventionally, neuroradiologists analyze DSA images visually for diagnosis, clinical decision-making, and postoperative evaluation, which is inherently time-consuming, labor-intensive, and subjective. Automated vessel segmentation in DSA promises to alleviate these challenges by accurately outlining vascular morphology. Currently, DSA-based cerebrovascular image segmentation finds utility in various downstream clinical applications, including automated thrombolysis in cerebral infarction scoring [5], 3D reconstruction of vessel [6], and image guidance of interventional surgery [7]. Despite its critical importance, cerebrovascular segmentation in DSA has not been extensively researched. A key factor contributing to this is the lack of a dedicated DSA dataset specifically designed for developing and evaluating segmentation techniques. However, the creation of such a dataset is particularly challenging due to the unique characteristics of DSA imaging. Primarily, in the DSA sequence, the intravascular contrast agent is projected from a fixed angle, revealing only a portion of the intracranial arteries within the contrast medium in each image, as illustrated in Fig. 1. Annotating vessels in DSA necessitates neuroradiologists to meticulously examine each image frame, a process that is both labor-intensive and costly [8]. Moreover, the intricate nature of cerebral vessels, characterized by overlapping structures, reduced contrast, and complex background noise, significantly increases the difficulty of accurate annotation.
@@
Currently, several segmentation approaches for DSA cerebral vessels have emerged [9, 10, 8]. However, these studies rely on in-house datasets, lacking a unified standard for evaluation. Most of these menthods [9, 10] focus on segmenting single-frame images, thereby ignoring DSA sequences’ inherent temporal attributes. As a result, single-frame based segmentation approaches may not fully represent the intricate vascular morphology of IA, potentially compromising their effectiveness in the diagnosis and treatment of cerebrovascular diseases [8]. Vessel segmentation in DSA is considered a dimensionality reduction task, where the input is 2D+Time images, and the output is a 2D segmentation map. This approach deviates from the traditional 2D and 3D vessel segmentation methods [11, 12, 13]. The key challenge lies in accurately capturing the spatiotemporal information present within the angiography for precise vessel segmentation.
@@
In this work, we propose the DSA-sequence Intracranial Artery Segmentation dataset (DIAS), meticulously annotated at the pixel level by neurosurgeons. Acknowledging the complexity and labor-intensiveness of vessel annotation in DSA sequences, we’ve provided two types of weak annotations for training samples to support research in weakly-supervised IA segmentation. To comprehensively evaluate the efficacy of prevalent segmentation methods like fully-supervised, weakly-supervised, and semi-supervised segmentation on the DIAS dataset, we have established a benchmark comprising a vessel sequence segmentation network, a scribble-based learning segmentation framework, and a random patch-based self-training semi-supervised segmentation pipeline. Extensive experiments with the DIAS set robust baselines, potentially advancing future algorithmic developments in cerebral vessel segmentation.
TABLE I: Summary of several public vessel segmentation datasets. ${}^{\dagger}$ represents only training set data is available.
@@
${}^{\dagger}$ [21] 2022 coronary artery CTA 3D 40 512$\times$512$\times$200
${}^{\dagger}$ [22] 2014 lung vessel CT 3D 20 -
@@ II-B Vessel Segmentation Methods
However, there have been few studies on DSA vessel segmentation. Meng et al. [9] proposed a CNN-based multiscale dense CNN segmentation network for IA segmentation in single-frame DSA, utilizing multiscale atrous convolution, enhanced dense blocks, and redesigned skip connections. Zhang et al. [23] proposed a boundary enhancement and feature denoising module to improve the extraction of boundary information in DSA coronary segmentation. These two approaches are tailored for vessel segmentation in individual DSA frames. For vessel sequence, Hao et al. proposed the sequential vessel segmentation network (SVS-Net) for vessel segmentation in X-ray coronary angiography, which incorporates 3D convolutional layers to extract rich temporal-spatial features and employs channel attention blocks to learn discriminative features [38]. Su et al. proposed a spatio-temporal U-Net (ST U-Net) architecture which integrates convolutional gated recurrent units (ConvGRU) in the contracting branch of U-Net for cerebral artery and vein segmentation in DSA sequences [8]. While both methods achieve vessel segmentation within sequences, the former is not adaptable to variable sequences, and the latter is challenging to extend to other models.
@@
Fully annotating vessel datasets is both time-consuming and costly. Consequently, annotation-efficient segmentation methods based on weakly-supervised and semi-supervised have attracted significant attention from researchers. Weakly-supervised segmentation (WSS) trains models with sparse annotations instead of full ones. The common methods involves online labeling [39] via consistency checks with the mean teacher or using pseudo-labels supervision [40] derived from model predictions. Semi-supervised segmentation (SSS) aims to explore tremendous unlabeled data with supervision from limited labeled data. Recently, self-training methods [41, 42] have dominated this field. Furthermore, methods employing consistency regularization strategies [41, 43, 44] improve the generalization ability by encouraging high similarity in predictions from two perturbed networks for the same input image. To our knowledge, there is no existing literature on semi-supervised or weakly-supervised vessel segmentation in DSA sequences. Building upon our created dataset, we have conducted in-depth research and proposed innovative methods in three distinct areas: network models, WSS, and SSS. These approaches are specifically tailored to leverage the unique features of the dataset, addressing each segmentation paradigm comprehensively.
@@ III-A Dataset Description
The DIAS dataset is derived from DSA images obtained during ischemic stroke treatments at Beijing Tiantan Hospital. This dataset encompasses cases of ICAS and MCAO. The ethical committee at Beijing Tiantan Hospital authorized the use of all the data presented in this study. All images were retrospectively collected between January 2019 and December 2021 and then anonymized to protect the patients’ privacy by removing all identifiable clinical treatment details. The original DSA sequences were captured by the Siemens angiography machine during the neurointerventional procedures and stored as DICOM files. It contains 1027 2D-DSA sequences from 416 patients in anteroposterior view or lateral view. In this study, we retained 120 sequences (753 images) after excluding those with significant motion artifacts or incomplete arterial phase, or any duplicates. A total of 120 sequences were derived from 60 patients, among whom 24 were female and 36 were male. The average age of these patients was 62.5 years, with an age range from 40 to 88 years old. The number of frames varies between sequences (6—46 frames), but the frame rate is fixed (4 FPS). The sequences were then extracted into individual images with a resolution of 800$\times$800 pixels using RadiAnt. We retained all the arterial-phase frames and removed the pre-contrast, capillary-phase, and venous-phase frames under the guidance of a neurosurgeon. As a result, the length of DSA sequences in the DIAS dataset ranges from 4 to 14 frames.
@@ III-B1 Full Annotation
Figure 2: Full annotated process of DIAS.
The manual annotation of IA in DSA sequences necessitates a thorough consideration of all image data within the sequence. Fig. 1 illustrates a typical representation of intracranial vessels across five successive frames during the arterial phase in a lateral view. During this phase, the contrast agent progressively spreads from the internal carotid artery (ICA) into cerebral vessels like the anterior cerebral artery (ACA) middle cerebral artery (MCA), carried by the blood flow. For a single image, the contrast agent is discernible in only a portion of the IA. For instance, the contrast agent at t filled the C1-C3 segment of the ICA, while at t+3, it filled the ACA and MCA. To annotate the IA effectively, each image in the sequence must be individually annotated, and then these annotations should be integrated to leverage the full vessel information available in the sequence, thereby ensuring precision in annotation. However, the challenge arises with the fine nature of vessels, especially in images from later arterial stages, which makes the standard approach of demarcating their borders impractical for medical image annotation. Thus, we have to annotate each vessel pixel by pixel, which is a labor-intensive and time-consuming procedure. In reality, the angiographic representations of vessels in adjacent frame images exhibit a high degree of similarity, indicating that annotating every pixel in each frame image involves a substantial amount of repetitive work. Fig. 2 presents our proposed solution, a pragmatic compromise that aims to reduce the workload without compromising the accuracy of annotations:
@@ III-B2 Scribble Annotation
Scribble-based image segmentation has proven to be an efficacious and sought-after method for WSS [45]. This approach to weak annotation allows annotators to easily mark object centers without the challenging task of delineating class boundaries. To explore the potential of scribble supervision for IA segmentation in DSA sequences, we generated two forms of scribble annotations for the training set of DIAS: the Scribbles Annotation with Little Clinical Experience (SALE) and Random Drawing based on the Full Annotation (RDFA). SALE simulates vessel annotation with little expert experience, wherein the thick vessels highlighted by the contrast agent are annotated, as illustrated by the red marker in Fig. 4 (a). Despite receiving clinicians’ basic training, accurately distinguishing the thin vessels and edges of thick blood vessels with low contrast remains challenging. Consequently, pixels in these regions were not annotated during annotation. Moreover, scribble segmentation relies on supervised information for every category. Apart from annotating the vessels, we conducted extensive scribble annotations for the background wherever possible, as illustrated by the blue marker in Fig. 4 (a). Follow [46], we generate RDFA by skeletonizing and randomly pruning each category of the full annotation, as shown in Fig. 4 (b). RDFA simulates random scribble annotations, diverging from SALE, which focuses on easier annotations. In SALE and RDFA, we set the pixel values of the background and vessel annotations to 0 and 1, respectively, while pixels without annotations are assigned a value of 255.

### v4

@@ Abstract
The automated segmentation of Intracranial Arteries (IA) in Digital Subtraction Angiography (DSA) plays a crucial role in the quantification of vascular morphology, significantly contributing to computer-assisted stroke research and clinical practice. Current research primarily focuses on the segmentation of single-frame DSA using proprietary datasets. However, these methods face challenges due to the inherent limitation of single-frame DSA, which only partially displays vascular contrast, thereby hindering accurate vascular structure representation. In this work, we introduce DIAS, a dataset specifically developed for IA segmentation in DSA sequences. We establish a comprehensive benchmark for evaluating DIAS, covering full, weak, and semi-supervised segmentation methods. Specifically, we propose the vessel sequence segmentation network, in which the sequence feature extraction module effectively captures spatiotemporal representations of intravascular contrast, achieving intracranial artery segmentation in 2D+Time DSA sequences. For weakly-supervised IA segmentation, we propose a novel scribble learning-based image segmentation framework, which, under the guidance of scribble labels, employs cross pseudo-supervision and consistency regularization to improve the performance of the segmentation network. Furthermore, we introduce the random patch-based self-training framework, aimed at alleviating the performance constraints encountered in IA segmentation due to the limited availability of annotated DSA data. Our extensive experiments on the DIAS dataset demonstrate the effectiveness of these methods as potential baselines for future research and clinical applications. The dataset and code are publicly available at https://doi.org/10.5281/zenodo.11396520 and https://github.com/lseventeen/DIAS.
@@ I Introduction
Cerebrovascular diseases substantially contribute to global mortality and long-term disability [1]. Recent decades have witnessed a decline in mortality from Intracranial Artery (IA) diseases like Intracranial Arterial Stenosis (ICAS), Middle Cerebral Artery Occlusion (MCAO), and intracranial aneurysm, thanks to advancements in medical and imaging technology [2]. Among these, Digital Subtraction Angiography (DSA), an imaging modality, visualizes the flow of a contrast agent through vessels by capturing a sequence of images over a duration typically spanning 3 to 15 seconds, with a sampling rate between 3 to 5 frames per second [3]. It effectively removes bony structures by subtracting pre-contrast images, thereby enhancing the visibility of vessels filled with the contrast agent. Owing to its inherently superior spatial and temporal resolution, DSA can accurately reveal lesion details in instances where Computed Tomography Angiography (CTA) and Magnetic Resonance Angiography (MRA) may not offer a definitive diagnosis [2]. Thus, DSA is universally acknowledged as the gold standard for investigating lesion angioarchitecture, deciphering the dynamics of arterial blood supply, and guiding endovascular intervention [4].
@@
Conventionally, neuroradiologists analyze DSA images visually for diagnosis, clinical decision-making, and postoperative evaluation, which is inherently time-consuming, labor-intensive, and subjective. Automated vessel segmentation in DSA promises to alleviate these challenges by accurately outlining vascular morphology. Currently, DSA-based cerebrovascular image segmentation finds utility in various downstream clinical applications, including automated thrombolysis in cerebral infarction scoring [5], 3D reconstruction of vessel [6], and image guidance of interventional surgery [7]. Despite its critical importance, cerebrovascular segmentation in DSA has not been extensively researched. A contributing factor to this issue is the lack of a dedicated dataset designed for the development and evaluation of vessel segmentation algorithms in DSA. However, the creation of such a dataset is particularly challenging due to the unique characteristics of DSA imaging. Primarily, in the DSA sequence, the intravascular contrast agent is projected from a fixed angle, revealing only a portion of the intracranial arteries within the contrast medium in each image, as illustrated in Fig. 1. Annotating vessels in DSA necessitates neuroradiologists to meticulously examine each image frame, a process that is both labor-intensive and costly [8]. Moreover, the intricate nature of cerebral vessels, characterized by overlapping structures, reduced contrast in fine vessels, and complex background noise, significantly increases the difficulty of accurate annotation.
@@
Currently, several studies [9, 10, 8, 11] have emerged in the domain of vessel segmentation within DSA, demonstrating exceptional performance on in-house datasets. Nevertheless, the majority of these methods [9, 10, 11] focus on segmenting single-frame images, thereby neglecting the inherent temporal attributes of DSA sequences. Constrained by the fact that single-frame DSA captures only a fraction of the vascular contrast, such methods may inadequately represent the complex vascular structure of IA. This shortfall could potentially undermine their utility in the accurate diagnosis and treatment of cerebrovascular diseases [8]. Vessel segmentation in DSA should be perceived as a dimensionality reduction task, where the input is 2D+Time images, and the output is a 2D segmentation map. This approach deviates from the traditional 2D and 3D vessel segmentation methods [12, 13, 14]. The key challenge lies in accurately capturing the spatiotemporal information present within the angiography for precise vessel segmentation.
@@
TABLE I: Summary of several public vessel segmentation datasets. represents only training set data is available.
@@
ASOCA [22] 2022 coronary artery CTA 3D 40 512$\times$512$\times$200
VESSEL12 [23] 2014 lung vessel CT 3D 20 -
@@
In this work, we propose the DSA-sequence Intracranial Artery Segmentation dataset (DIAS), meticulously annotated at the pixel level by neurosurgeons. Acknowledging the complexity and labor-intensiveness of vessel annotation in DSA sequences, we’ve provided two types of weak annotations for training samples to support research in weakly-supervised IA segmentation. To comprehensively evaluate the efficacy of prevalent segmentation methods, including fully-supervised, weakly-supervised, and semi-supervised techniques, on the DIAS dataset, we have established a benchmark comprising three key components:
  • We propose the vessel sequence segmentation network (VSS-Net), incorporating the sequence feature extraction module adept at efficiently capturing spatiotemporal representations of intravascular contrast. Notably, VSS-Net is designed to be compatible with any encoder-decoder network, enabling precise segmentation of intracranial arteries in 2D+Time DSA sequences.
  • We propose an innovative scribble learning-based image segmentation framework, specifically designed for weakly-supervised IA segmentation. This framework, distinguished by its dual-model architecture, employs cross pseudo-supervision and consistency regularization, guided by scribble labels. This integration significantly boosts the performance of the segmentation network, capitalizing on the synergistic advantages offered by the dual-model approach.
  • We introduce the random patch-based self-training algorithm, which alleviates the challenges arising from the limited availability of annotated DSA data in IA segmentation by enhancing the model’s generalization through joint training on randomly selected image patches from both labeled and pseudo-labeled data.
Through extensive experimentation on the DIAS dataset, we have established robust baselines that hold the potential to advance future algorithmic developments in cerebral vessel segmentation.
@@ II-B Vessel Segmentation Methods
There have been few studies on DSA vessel segmentation. [9] proposed a CNN-based multiscale dense CNN segmentation network for IA segmentation in single-frame DSA, utilizing multiscale atrous convolution, enhanced dense blocks, and redesigned skip connections. [24] proposed a boundary enhancement and feature denoising module to improve the extraction of boundary information in DSA coronary segmentation. These two approaches are tailored for vessel segmentation in individual DSA frames. For vessel sequence, [39] proposed the sequential vessel segmentation network (SVS-Net) for vessel segmentation in X-ray coronary angiography, which incorporates 3D convolutional layers to extract rich temporal-spatial features and employs channel attention blocks to learn discriminative features. [8] proposed a spatio-temporal U-Net (ST U-Net) architecture which integrates convolutional gated recurrent units (ConvGRU) in the contracting branch of U-Net for cerebral artery and vein segmentation in DSA sequences. While both methods achieve vessel segmentation within sequences, the former is not adaptable to variable sequences, and the latter is challenging to extend to other models.
@@
Fully annotating vessel datasets is both time-consuming and costly. Consequently, annotation-efficient segmentation methods based on weakly-supervised and semi-supervised have attracted significant attention from researchers. Weakly-Supervised Segmentation (WSS) trains models with sparse or noisy annotations, instead of exact and complete annotations. The common methods involves online labeling [40] via consistency checks with the mean teacher or using pseudo-labels supervision [41] derived from model predictions. Furthermore, investigating the method for weakly-supervised annotation of vessels in DSA is a research direction deserving of exploration. [11] propose a novel learning framework, which utilizes an active contour model for weak supervision and low-cost human-in-the-loop strategies to improve weak label quality. Semi-Supervised Segmentation (SSS) aims to explore tremendous unlabeled data with supervision from limited labeled data. Recently, self-training methods [42, 43] have dominated this field. Furthermore, methods employing consistency regularization strategies [42, 44, 45] improve the generalization ability by encouraging high similarity in predictions from two perturbed networks for the same input image.
Various domain adaptation techniques have been proposed to enhance the generalizability of vessel segmentation, leveraging vessel data from other modalities with similar anatomical structures to aid in training, thereby improving the performance of models in SSS. [46] propose contrastive SSS for cross anatomy domain adaptation that adapts a model to segment coronary arteries, which requires only limited annotations in the coronary arteries by leveraging existing annotated retinal vessels of similar structure. [47] propose a semi-supervised domain adaptation framework for accurately segmenting 3D brain vessels by employing the StyleGAN2 [48] to facilitate the representation of heterogeneous volumetric data and bridge the substantial domain gap between angiography and venography brain images. To our knowledge, there is no existing literature on weakly-supervised and semi-supervised vessel segmentation in DSA sequences. Building upon our created dataset, we have conducted in-depth research and proposed innovative methods in three distinct areas: network models, WSS, and SSS. These approaches are specifically tailored to leverage the unique features of the dataset, addressing each segmentation paradigm comprehensively.
Furthermore, large-scale pretrained models, i.e., foundation models, exhibit considerable generalizability and adaptability, positioning them as a focal point of recent medical image analysis research [49]. However, generally, a unified foundation model approach cannot achieve state-of-the-art performance in many medical image analysis tasks due to large variations present in organs and important structures, texture, shape, size and topology (e.g., blood vessels), and imaging modalities [50]. We are convinced that, through continuous research, foundation models possess the potential to achieve vascular segmentation in DSA that can compete with fully supervised methods, leveraging few-shot learning, zero-shot learning, or prompt engineering.
@@ III-A Dataset Description
The DIAS dataset is derived from DSA images obtained during ischemic stroke treatments at Beijing Tiantan Hospital. This dataset encompasses cases of ICAS and MCAO. The ethical committee at Beijing Tiantan Hospital authorized the use of all the data presented in this study. All images were retrospectively collected between January 2019 and December 2021 and then anonymized to protect the patients’ privacy by removing all identifiable clinical treatment details. The original DSA sequences were captured by the Siemens angiography machine during the neurointerventional procedures and stored as DICOM files. It contains 1027 2D-DSA sequences from 416 patients in anteroposterior view or lateral view. In this study, we retained 120 sequences (762 images) after excluding those with significant motion artifacts or incomplete arterial phase, or any duplicates. A total of 120 sequences were derived from 60 patients, among whom 24 were female and 36 were male. The average age of these patients was 62.5 years, with an age range from 40 to 88 years old. The number of frames varies between sequences (6 - 46 frames), but the frame rate is fixed (4 FPS). The sequences were then extracted into individual images with a resolution of 800$\times$800 pixels using RadiAnt. We retained all the arterial-phase frames and removed the pre-contrast, capillary-phase, and venous-phase frames under the guidance of a neurosurgeon. As a result, the length of DSA sequences in the DIAS dataset ranges from 4 to 14 frames.
Figure 2: Full annotated process of DIAS.
@@ III-B1 Full Annotation
The manual annotation of IA in DSA sequences necessitates a thorough consideration of all image data within the sequence. Fig. 1 illustrates a typical representation of intracranial vessels across five successive frames during the arterial phase in a lateral view. During this phase, the contrast agent progressively spreads from the Internal Carotid Artery (ICA) into cerebral vessels like the Anterior Cerebral Artery (ACA) and Middle Cerebral Artery (MCA), carried by the blood flow. For a single image, the contrast agent is discernible in only a portion of the IA. For instance, the contrast agent at t filled the C1-C3 segment of the ICA, while at t+5, it filled the ACA and MCA. To annotate the IA effectively, each image in the sequence must be individually annotated, and then these annotations should be integrated to leverage the full vessel information available in the sequence, thereby ensuring precision in annotation. However, the challenge arises with the fine nature of vessels, especially in images from later arterial stages, which makes the standard approach of demarcating their borders impractical for medical image annotation. Thus, we have to annotate each vessel pixel by pixel, which is a labor-intensive and time-consuming procedure. In reality, the angiographic representations of vessels in adjacent frame images exhibit a high degree of similarity, indicating that annotating every pixel in each frame image involves a substantial amount of repetitive work. Fig. 2 presents our proposed solution, a pragmatic compromise that aims to reduce the workload without compromising the accuracy of annotations:
@@ III-B2 Scribble Annotation
Scribble-based image segmentation has proven to be an efficacious and sought-after method for WSS [51]. This approach to weak annotation allows annotators to easily mark object centers without the challenging task of delineating class boundaries. To explore the potential of scribble supervision for IA segmentation in DSA sequences, we generated two forms of scribble annotations for the training set of DIAS: the Scribbles Annotation with Little Clinical Experience (SALE) and Random Drawing based on the Full Annotation (RDFA). SALE simulates vessel annotation with little expert experience, wherein the thick vessels highlighted by the contrast agent are annotated, as illustrated by the red marker in Fig. 4 (a). Despite receiving clinicians’ basic training, accurately distinguishing the thin vessels and edges of thick blood vessels with low contrast remains challenging. Consequently, pixels in these regions were not annotated during annotation. Moreover, scribble segmentation relies on supervised information for every category. Apart from annotating the vessels, we conducted extensive scribble annotations for the background wherever possible, as illustrated by the blue marker in Fig. 4 (a). Follow [52], we generate RDFA by skeletonizing and randomly pruning each category of the full annotation, as shown in Fig. 4 (b). RDFA simulates random scribble annotations, diverging from SALE, which focuses on easier annotations. In SALE and RDFA, we set the pixel values of the background annotation and vessel annotation to 0 and 1, respectively, while pixels without annotations are assigned a value of 255.